OpenStudy (anonymous):

Fe2O3 + 3CO --> 2Fe +3CO2 determine the moles of CO required to make 10g of iron.

OpenStudy (nottim):

Could you redo that using the equation button?

OpenStudy (nottim):

I wanna know which ones are subscripts.

OpenStudy (nottim):

Hey, you still alive?

OpenStudy (nottim):

Also, you know the steps to reach the solution right?

OpenStudy (nottim):

First, convert the mass of the iron to moles. then, find the ratio between the 2. Once you're done that, find the moles of CO.

OpenStudy (anonymous):

oh right yeah im still alive just exploded myself :(

OpenStudy (anonymous):

lol kay soo umm anyone know how to do this question

OpenStudy (anonymous):

the ans is (10/56)*(3/2)

OpenStudy (anonymous):

15/56

OpenStudy (anonymous):

moles

OpenStudy (anonymous):

but the answer is 5 moles thats what im confused about i got the same answer as you

OpenStudy (anonymous):

indian??

OpenStudy (anonymous):

are u talking about my teacher or me :P

OpenStudy (anonymous):

u...

OpenStudy (anonymous):

ok so the answer that we got is right then?

OpenStudy (anonymous):

which book is this ques frm??

OpenStudy (anonymous):

i am 99% confident about this ans...

OpenStudy (anonymous):

well its from an online document my teacher posted and yeah same here

OpenStudy (anonymous):

ask ur teacher then...

OpenStudy (anonymous):

alri thanks

OpenStudy (anonymous):

\[10.0 g Fe (\frac{1 mole Fe}{55.85 g Fe})(\frac{3 mole CO}{2 mole Fe})(\frac{28.01 g CO}{1 mole CO})=7.52 g CO\]

OpenStudy (anonymous):

\[10.0 g Fe (\frac{1 mole Fe}{55.85 g Fe})(\frac{3 mole CO}{2 mole Fe})=0.269 mole CO\] I went to grams with the first one then read you needed moles. So i redid it and you need .269 moles of CO to make 10 g Fe.

OpenStudy (anonymous):

thanks :)